Find an equation of the ellipse satisfying the given conditions.
Center at (0,0), one endpoint of the major axis at (6,0), and one endpoint of the minor axis at (0,4)

Answer by BryceWalker(4)

The center is (0,0)
The major axis has an endpoint of (6,0) and the other one is (x,y)
The minor axis has an endpoint of (0,4) and the other one is (x,y)
The major axis of this would be x=0 and the minor axis would be y=0
Half of the length of the major axis is 6 multiplied by 2 ,; length would be 12
Half of the length of the minor axis is 4 multiplied by 2 ,; length would be 8
To find a: the formula for the length of M.A when the ellipse is vertical is 2a; therefore a would be 6
To find be: its the same but the formula would be 2b ; therefore b would 4
a^2 = 36
b^2 = 16
Then, the formula would be x^2/16 + y^2/36 = 1
(^) means raise to the power of
(/) means division: divided by
I don't know if this is the correct procedure. Hope it solves the question